### v1

# Neural Diversity Regularizes Hallucinations in Small Language Models

Abstract
Language models continue to hallucinate despite increases in parameters, compute, and data. We propose neural diversity — decorrelated parallel representations — as a principled mechanism that reduces hallucination rates at fixed parameter and data budgets. Inspired by portfolio theory, where uncorrelated assets reduce risk by $\sqrt{P}$, we prove hallucination probability is bounded by representational correlation: $\mathbb{P}(\mathrm{H})\leq f(\sigma^{2}((1-\rho(P))/P+\rho(P)),\mu^{2})$, which predicts that language models need an optimal amount of neurodiversity. To validate this, we introduce ND-LoRA (Neural Diversity Low-Rank Adaptation), combining parallel LoRA adapters with Barlow Twins regularization, and demonstrate that ND-LoRA reduces hallucinations by up to 25.6% (14.6% avg.) without degrading general accuracy. Ablations show LoRA adapters and regularization act synergistically, causal interventions prove neurodiversity as the mediating factor and correlational analyses indicate scale: a 0.1% neural correlation increase is associated with a 3.8% hallucination increase. Finally, task-dependent optimality emerges: different tasks require different amounts of optimal neurodiversity. Together, our results highlight neural diversity as a third axis of scaling — orthogonal to parameters and data — to improve the reliability of language models at fixed budgets.
Figure 1: Maximizing reliability requires an optimal amount of neural diversity. We vary parallel decorrelated representations ($P\in\{1,2,4,8\}$) and measure performance across 6 hallucination benchmarks. The x-axis shows distance from the performance-maximizing $P_{\star}$: $\Delta P=P-P_{\star}$. Aggregating 182,850 evaluation points (LOWESS fit, 80% bootstrap CI, 1000 bootstrap resamplings), the U-shaped curve shows performance improves rapidly with neural diversity, peaking at a specific neurodiversity optimum and then degrading slowly with excessive diversity.
1 Introduction
Despite scaling to trillions of parameters, language models hallucinate persistently (Lin et al., 2021). This reliability crisis is acute for small language models (SLMs) — increasingly favored for edge deployment (Chen et al., 2024) — whose compressed representations make them especially vulnerable to hallucinations, with even well-resourced efforts like GPT-OSS 20B exhibiting 91% hallucination rates on factual benchmarks (OpenAI, 2025).
Existing scaling methods — parameters, inference-time compute, and parallel architectures like ParScale (Chen et al., 2025) — predominantly target accuracy (first-moment improvements in perplexity and task performance). Yet reliability (second-moment reduction in hallucinations and factual errors) remains largely unaddressed. While ParScale achieves $O(\log P)$ performance gains with better memory and latency, it provides no explicit mechanism for reducing hallucinations. Without explicit regularization, representational collapse (Jing et al., 2022) drives parallel streams toward identical features, leaving reliability gains unrealized despite added computational resources.
Drawing from portfolio theory, we propose that decorrelated representations reduce hallucination risk just as uncorrelated assets reduce financial risk. When streams are perfectly correlated ($\rho=1$), parallelism provides no benefit; when decorrelated ($\rho\to 0$), signal-to-noise improves by $\sqrt{P}$.
We formalize this through a neural diversity framework proving hallucination probability is bounded by cross-stream correlation: $\mathbb{P}(\text{hallucination})\leq f(\sigma^{2}((1-\rho)/P+\rho),\mu^{2})$. This reveals why existing scaling approaches don’t address reliability and how explicit decorrelation provides a mechanism to reduce hallucinations at fixed parameter budgets.
To validate this, we introduce ND-LoRA, combining independent LoRA adapters with Barlow Twins regularization (Zbontar et al., 2021). Despite adding only 0.004% more computation, ND-LoRA reduces hallucinations by up to 25.6% (14.6% avg.) while preserving general performance.
Our experiments establish: (1) neural diversity causally reduces hallucination probability; (2) there is an optimal amount of neural diversity and that optimal amount is task-dependent; and (3) the relationship between neural diversity and hallucination probability is theoretically provable.
Neural diversity represents a third scaling axis beyond parameters and data. While traditional scaling asks “how big?” and data scaling “how much?”, diversity scaling asks “how different?” — crucial for achieving reliability without massive computational investment.
Our contributions are:
  • Theory: The first principled connection between neural spectral diversity and hallucination probability that matches empirically observed hallucination rates
[…]
  • Analysis: Evidence of the concavity of $P$ in industry-standard benchmarks and the presence of task-dependent optimal $P$ revealing when neural diversity helps most (and hurts)
  • Framework: A general principle for parallel architectures — that maintaining decorrelation is as important as adding capacity — with implications beyond hallucination mitigation
2 A Theory of Neural Diversity
Why don’t existing scaling methods improve reliability? Without explicit diversity mechanisms, gradient descent drives parallel streams toward similar representations through representational collapse (Jing et al., 2022), leaving reliability gains unrealized. We establish the first rigorous connection between architectural diversity and hallucination probability, proving that decorrelated streams reduce hallucinations and providing mathematical foundations for ND-LoRA.
2.1 Preliminaries
We develop a framework connecting parallel neural computation to hallucination probability. Language models produce unreliable outputs in several ways: fabricating facts contradicting known knowledge, generating content inconsistent with input context, or creating unsupported claims (Maynez et al., 2020; Ji et al., 2023). While recent surveys provide comprehensive taxonomies (Huang et al., 2024; Tonmoy et al., 2024), we model hallucination probability through a signal-noise analytical proxy that captures the underlying reliability failure common to these manifestations and is analytically tractable.
Consider a language model processing input $x$ to produce an output (e.g., answer to a question, summary of a document). For each input, assume an oracle truth $T>0$ exists. Our architecture employs $P$ parallel computational pathways called streams, each processing the same input through the same model but in perturbed ways. Each stream $i$ produces output $\hat{Y}_{i}=T+m_{i}$, where $m_{i}$ represents zero-mean noise corrupting the truth. These streams are combined through a weighted average:
$M\;\triangleq\;\frac{1}{P}\sum_{i=1}^{P}\hat{Y}_{i}\;=\;T+\frac{1}{P}\sum_{i=1}^{P}m_{i}.$ (1)
We assume $M$ has learned an unbiased truth, so $\mathbb{E}[M]=T\equiv\mu>0$. Each stream’s noise has mean $\mathbb{E}[m_{i}]=0$, variance $\mathrm{Var}(m_{i})=\sigma^{2}$, and pairwise correlations $\rho_{ij}=\mathrm{Corr}(m_{i},m_{j})$ measuring representational collapse. The average pairwise correlation captures overall similarity:
$\rho\;\triangleq\;\frac{2}{P(P-1)}\sum_{1\leq i<j\leq P}\rho_{ij}.$ (2)
We define the hallucination event as $\mathrm{H}\equiv\{M\leq 0\}$—when aggregated noise overwhelms the true signal, causing incorrect predictions.
The key intuition is noise cancellation: when streams are perfectly correlated ($\rho=1$), noise doesn’t average out and $M$ remains noisy. When decorrelated ($\rho\to 0$), independent noise cancels and $M$ concentrates around truth $T$, reducing hallucination probability by $\sqrt{P}$.
2.2 Neural Diversity Bounds Hallucination
We connect noise correlation to feature correlation through linearization: if noise terms depend approximately linearly on representations at a chosen design layer, then reducing feature correlation reduces noise correlation.
Design-layer features and diversity. At a chosen design layer, stream $i$ exposes a feature vector $z_{i}\in\mathbb{R}^{d}$. Let $\tilde{z}_{i}$ denote a whitened version (zero-mean, identity covariance within stream), and define the cross-correlation matrices $C_{ij}\triangleq\mathbb{E}[\tilde{z}_{i}\tilde{z}_{j}^{\top}]$.
Definition 1 (Neural Diversity Index).
The spectral diversity index is
$\mathcal{D}_{\mathrm{spec}}\;=\;\frac{1}{P(P-1)}\sum_{i\neq j}\,\left\lVert C_{ij}\right\rVert_{2}.$ (3)
Lower $\mathcal{D}_{\mathrm{spec}}$ indicates greater diversity: $\mathcal{D}_{\mathrm{spec}}=0$ means perfectly decorrelated streams, while $\mathcal{D}_{\mathrm{spec}}=1$ means complete collapse.
We now connect stream correlation to hallucination probability through three steps: variance analysis, linearization, and the main bound.
Lemma 1 (Variance of Aggregated Output).
Under the assumptions above,
$\mathrm{Var}(M)\;=\;\sigma^{2}\!\left(\frac{1-\rho}{P}+\rho\right).$ (4)
Proof sketch.
Expand $\mathrm{Var}(M)$ via $\mathrm{Var}(\frac{1}{P}\sum m_{i})$, use $\mathrm{Var}(m_{i})=\sigma^{2}$ and $\mathrm{Cov}(m_{i},m_{j})=\rho\,\sigma^{2}$ for $i\neq j$, then collect terms.
When $\rho=0$, variance decreases as $\sigma^{2}/P$. When $\rho=1$, variance stays at $\sigma^{2}$ regardless of $P$.
Lemma 2 (Correlation Bound).
Assume the noise terms are locally linear in whitened features, $m_{i}=v_{i}^{\top}\tilde{z}_{i}$, and let $\sigma_{i}^{2}\triangleq\mathrm{Var}(m_{i})$. Then for $i\neq j$,
$\rho_{ij}\;=\;\frac{\mathrm{Cov}(m_{i},m_{j})}{\sigma_{i}\sigma_{j}}\;\leq\;\kappa_{ij}\,\left\lVert C_{ij}\right\rVert_{2},\qquad\kappa_{ij}\;\triangleq\;\frac{\left\lVert v_{i}\right\rVert\,\left\lVert v_{j}\right\rVert}{\sigma_{i}\sigma_{j}}.$ (5)
Proof sketch.
Compute $\mathrm{Cov}(m_{i},m_{j})=v_{i}^{\top}C_{ij}v_{j}$ and bound by the spectral norm: $|v_{i}^{\top}C_{ij}v_{j}|\leq\left\lVert v_{i}\right\rVert\,\left\lVert C_{ij}\right\rVert_{2}\,\left\lVert v_{j}\right\rVert$. Divide by $\sigma_{i}\sigma_{j}$.
Noise correlation $\rho_{ij}$ is controlled by feature correlation $\|C_{ij}\|_{2}$. We now have a direct path from $\mathcal{D}_{\mathrm{spec}}$ to hallucination probability.
Theorem 1 (Hallucination Probability Bound with Diversity).
The hallucination probability satisfies
$\mathbb{P}(\mathrm{H})\;\leq\;\frac{\sigma^{2}\left(\tfrac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right)}{\sigma^{2}\left(\tfrac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right)+\mu^{2}}\;+\;h_{0},$ (6)
where $\bar{\kappa}$ is the average correlation scaling factor and $0\leq h_{0}\leq 1$ is a constant accounting for approximation error.
Proof sketch.
By Lemma 1, $\mathrm{Var}(M)$ is increasing in $\rho$ for $P>1$. $\bar{\rho}\leq\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}$, hence $\mathrm{Var}(M)\leq\sigma^{2}\!\big(\tfrac{1-\bar{\kappa}\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\mathcal{D}_{\mathrm{spec}}\big)$. Plug this variance bound into Cantelli’s bound. If the linear/whitening conditions fail on a set of probability $h_{0}$, upper bound that set trivially by 1 and add $h_{0}$.
This gets us the first half of our theoretical result: Lower $\mathcal{D}_{\mathrm{spec}}$ reduces hallucination probability and, additionally, the benefit scales with $P$.
2.3 Scaling Behavior
Next, we demonstrate that under common circumstances, the hallucination bound follows a U-shaped curve — initially decreasing with higher $P$, but starts increasing eventually. Consider the case where the correlation itself increases with $P$, say, due to optimizer constraints:
Theorem 2 (“U-shape” of the Hallucination Bound under Rising Correlation).
Suppose the average correlation increases with $P$ as $\rho(P)=\rho_{0}+\beta(P-1)^{\gamma}$ with $\beta>0$ and $\gamma>0$, and consider the Cantelli bound:
$\mathcal{B}(P)\;\triangleq\;\frac{\sigma^{2}\left(\tfrac{1-\rho(P)}{P}+\rho(P)\right)}{\sigma^{2}\left(\tfrac{1-\rho(P)}{P}+\rho(P)\right)+\mu^{2}}.$ (7)
Then $\mathcal{B}(P)$ is decreasing for $P$ near $1$ and increasing for $P$ sufficiently large. Moreover, if $\gamma\geq 1$ (so $\rho^{\prime\prime}(P)\geq 0$), $\mathcal{B}(P)$ is convex on $(1,\infty)$ and thus has a unique minimizer $P_{\star}>1$.
Proof sketch.
Work with the unnormalized variance factor $g(P)=\tfrac{1-\rho(P)}{P}+\rho(P)$. Differentiate: $g^{\prime}(P)=\rho^{\prime}(P)\big(1-\tfrac{1}{P}\big)-\tfrac{1-\rho(P)}{P^{2}}$. At $P\downarrow 1$, the first term vanishes while the second is negative, so $g^{\prime}(P)<0$. As $P\to\infty$, $\rho^{\prime}(P)>0$ dominates the $O(P^{-2})$ negative term, so $g^{\prime}(P)>0$ eventually. If $\gamma\geq 1$, then $g^{\prime\prime}(P)\geq 0$, giving a unique minimizer. Since $\mathcal{B}(P)$ is an increasing function of $g(P)$, the same qualitative behavior holds for $\mathcal{B}(P)$.
All together, Theorems 1 and 2 show that there exists an optimal $P_{\star}$ that minimizes hallucinations. Next, we show how to construct an architecture and training protocol to find $P_{\star}$.
3 ND-LoRA: A Practical Demonstration
Figure 2: ND-LoRA schematic for $P=4$ parallel streams. Each stream receives independent LoRA adapters and learnable prefix tokens. The aggregator combines stream outputs with learnable weights, while Barlow Twins regularization incentivizes decorrelation between stream outputs.
We introduce ND-LoRA (Neural Diversity Low-Rank Adaptation), a parameter-efficient method that demonstrates our theoretical framework for neural diversity regularization. ND-LoRA extends the ParScale architecture with stream-aware LoRA adapters and explicit decorrelation objectives.
3.1 Architecture
Our implementation builds on ParScale with $P$ parallel computation streams. Each stream $i\in\{1,\ldots,P\}$ uses 48 learnable prefix tokens prepended to the input sequence that flow through all layers via the attention mechanism, along with stream-specific LoRA adapters applied at each layer:
$h_{i}^{(\ell)}=\text{Layer}^{(\ell)}(h_{i}^{(\ell-1)}+B_{i}^{(\ell)}A_{i}^{(\ell)}h_{i}^{(\ell-1)})$ (8)
where $B_{i}^{(\ell)}\in\mathbb{R}^{d\times r}$, $A_{i}^{(\ell)}\in\mathbb{R}^{r\times d}$ are stream-specific LoRA matrices with rank $r$. The final output combines streams through a learned aggregator:
$y=\text{LM\_Head}\left(\sum_{i=1}^{P}\alpha_{i}\cdot h_{i}^{(L)}\right)$ (9)
where $\alpha_{i}=(1-\varepsilon)\cdot\text{softmax}(\text{MLP}([h_{1}^{(L)},\ldots,h_{P}^{(L)}]))_{i}+\varepsilon/P$ are dynamic weights with label smoothing ($\varepsilon=0.1$) computed from the concatenated stream representations. This prevents attention collapse by ensuring minimum weight $\varepsilon/P$ for each stream.
This architecture enables stream specialization while maintaining parameter efficiency. For $P=2$ streams with rank-16 LoRA, we use approximately 29K trainable parameters per layer, comparable to a single rank-32 LoRA but with fundamentally different representational capabilities.
3.2 Barlow Twins Regularization
To encourage neural diversity, we apply Barlow Twins regularization at a pre-specified design layer $\ell^{*}$. Let $z_{i}^{(\ell^{*})}\in\mathbb{R}^{B\times T\times d}$ denote the hidden representations of stream $i$ at the design layer. We first apply batch normalization and mean-centering to obtain whitened features $\tilde{z}_{i}$.
The cross-correlation matrix between streams $i$ and $j$ is:
$C_{ij}=\frac{1}{BT}\sum_{b,t}\tilde{z}_{i,bt}\tilde{z}_{j,bt}^{\top}\in\mathbb{R}^{d\times d}$ (10)
Our Barlow Twins loss promotes decorrelation by penalizing off-diagonal correlations:
$\mathcal{L}_{BT}=\frac{1}{P(P-1)}\sum_{i\neq j}\|C_{ij}-I\|_{F}^{2}$ (11)
The standard formulation scales quadratically with $P$, creating $\binom{P}{2}$ optimization dependencies that inhibit convergence. To address this scalability challenge, we implement a RandK variant that samples stream pairs stochastically:
$\mathcal{L}_{BT}=\mathbb{E}_{(i,j)\sim\text{MultN}(C_{i})}\|C_{ij}-I\|_{F}^{2}$ (12)
where $C_{i}$ induces a multinomial distribution over stream pairs. This reduces complexity from $O(P^{2})$ to $O(PK)$ where $K$ is the number of sampled pairs, enabling efficient scaling to larger $P$.
The total training objective combines cross-entropy and decorrelation terms:
$\mathcal{L}=\mathcal{L}_{CE}+\lambda_{BT}\mathcal{L}_{BT}$ (13)
4 Experimental Validation
4.1 Experimental Setup
We validate our theoretical framework through systematic hallucination reduction experiments using neural diversity regularization with parameter- and data-matched comparisons.
Model and Architecture. We use Qwen2.5-0.5B (896 hidden dimensions, 24 layers) with ND-LoRA across $P\in\{1,2,4,8\}$ parallel streams applied to QKV self-attention modules and a design layer of 20 for de-correlation loss. Each stream uses independent rank-16 LoRA adapters and 48 prefix tokens, totaling 5-20M trainable parameters with 494M backbone frozen. Baseline methods use higher-rank LoRA (R32-R128) for parameter matching.
Training Protocol. Models train on 20M tokens from The Pile (8 random shards, fixed seeds). We use 1024-token sequences, AdamW optimization (peak lr 3e-4, cosine decay, 2% warmup), batch size 64, bfloat16 precision, design layer 16. Training completes in $\sim$5K steps ($\sim$30 min. on A100).
Evaluation Benchmarks. We evaluate across: (1) Hallucination-sensitive: TruthfulQA (Lin et al., 2021), HaluEval (Li et al., 2023a), MemoTrap (McKenzie et al., 2023); (2) Knowledge-intensive: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023); (3) General capability: Wikitext BPB (Merity et al., 2017), Winogrande (Sakaguchi et al., 2020). This tests if neural diversity improves reliability without sacrificing general performance.
Statistical Methodology. We evaluate significance using McNemar’s test for binary classification tasks and two-tailed bootstrap tests with 10,000 samples for other tasks. Improvements marked with * are significant at $p<0.05$.
4.2 Key Results
Table 1 demonstrates ND-LoRA achieves substantial improvements on hallucination-sensitive benchmarks while maintaining competitive general performance. ND-LoRA with $P=2$ streams achieves statistically significant improvements on HaluEval-Summarization (0.481* vs 0.400, $p<0.001$, 8.1% absolute / 20.2% relative), TruthfulQA-MC2 (0.442* vs 0.403, $p=0.030$, 3.9% absolute / 9.5% relative) and MemoTrap (0.666* vs 0.634, $p<0.001$, 3.2% absolute / 5.1% relative) vs parameter-matched Qwen, validating our theoretical prediction.
Although ND-LoRA’s improvements specifically target reliability benchmarks, they preserve general capabilities. Qwen slightly outperforms on Wikitext (0.784 vs 0.778) and Natural Questions (0.055 vs. 0.065), but ND-LoRA wins on Winogrande (0.574). This task-selective pattern aligns with our framework for non-monotone behaviors and task-dependent optimality.
Parameter efficiency is evident comparing ND-LoRA R16 ($P=2$) against Qwen2.5-0.5B LoRA R32. Despite lower-rank adapters, ND-LoRA consistently outperforms the high-rank baseline on hallucination tasks, demonstrating that architectural diversity provides more value at equal capacity. This shows representational diversity, not parameter count, drives reliability gains in our experiment.
Model HaluEval MemoTrap TruthfulQA NQ Wikitext WG
ND-LoRA R16 (P=2) 0.481* 0.666* 0.442* 0.055 0.784 0.574
ParScale R32 (P=2) 0.439 0.638 0.412 0.059 0.793 0.564
Qwen LoRA R32 0.400 0.634 0.403 0.065 0.778 0.572
Table 1: Even at $P=2$ streams, ND-LoRA achieves up to 20.2% relative hallucination reduction vs. parameter-matched baseline. Across hallucination benchmarks, ND-LoRA shows statistically significant improvements (HaluEval-Summarization +20.2%*, MemoTrap +5.1%*, TruthfulQA-MC2 +9.5%*) while maintaining competitive Winogrande, NQ, and Wikitext general-purpose capabilities. Baselines use higher LoRA ranks for parameter parity. * indicates $p<0.05$.
Figure 3: Reliability improves as neural diversity increases (lower $\mathcal{D}_{\text{spec}}$). Specifically, spectral diversity ($\mathcal{D}_{\text{spec}}$) is negatively correlated with HaluEval-Summarization performance (slope=-37.842, R²=0.237, p=0.002), consistent with $\mathbb{P}(\text{H})\propto\mathcal{D}_{\text{spec}}$ in Theorem 1.
These findings establish neural diversity as a practical reliability mechanism. Consistent improvements across hallucination benchmarks with preserved general performance suggest ND-LoRA addresses fundamental reliability challenges rather than metric-specific optimization. Figure 3 demonstrates strong empirical correlation between neural diversity and performance, building intuition for the causal relationship established in Section 5.
4.3 Scaling Analysis
Our theoretical framework predicts U-shaped scaling curves where hallucination probability initially decreases with $P$ but increases as correlation rises without diversity regularization. Figure 1 empirically validates this across tasks, revealing a critical insight: maximizing reliability requires an optimal amount of neural diversity — more diversity is not always better.
4.4 Task-Dependent Optimality
Category Task Best $P$ Best Score $\Delta$}% Score Sig.
Hallucination HaluEval (Dialog) 4 0.516 +12.8% ***
HaluEval (QA) 4 0.451 +23.4% ***
HaluEval (Summ) 4 0.502 +25.6% ***
MemoTrap v2 8 0.689 +8.8% ***
TruthfulQA (MC1) 2 0.269 +7.3%
TruthfulQA (MC2) 2 0.442 +9.5% *
Knowledge NQ (8-shot) 1 0.066
NQ-swap 8 0.554 +0.8%
PopQA 1 0.111
TriviaQA (8-shot) 1 0.192
Table 2: The optimal amount of neural diversity varies by task. Hallucination tasks show large gains with optimal $P\in\{2,4,8\}$ (up to +25.6%), while knowledge tasks mostly peak at $P=1$. Even among hallucination tasks, optimal $P$ differs: HaluEval variants peak at $P=4$, TruthfulQA at $P=2$, MemoTrap at $P=8$. Each row shows the performance peak (best $P$), best score, relative increase vs. $P=1$ baseline, and statistical significance (*** for $p<0.001$, * for $p<0.05$).
Further, the optimal diversity is task-dependent. Table 2 reveals striking task-dependent sensitivity patterns relative to the $P=1$ baseline. Hallucination-focused tasks show the largest gains: HaluEval Summarization achieves +25.6% relative improvement at $P=4$, HaluEval QA shows +23.4% at $P=4$, and TruthfulQA MC2 shows +9.5% at $P=2$ while MemoTrap benefits from higher diversity ($P=8$, +8.8%). Notably, knowledge-intensive tasks like PopQA, TriviaQA and NQ show no improvement over baseline, which is expected as ND-LoRA does not add new sources of knowledge or try to improve recall of existing knowledge. This heterogeneity demonstrates that different tasks require different amounts of neural diversity to maximize reliability, with hallucination-focused tasks generally benefiting most from decorrelated representations.
5 Mechanistic Analysis
5.1 Neural Diversity as the Causal Mediator
To establish causality beyond correlation, we perform artificial corruption interventions that directly manipulate cross-stream similarity.
Experiment Design. Starting with a pre-trained ND-LoRA $P=4$ model, we inject a corruption hook at the RMSNorm layer that randomly substitutes the hidden state at randomly-chosen positions in a given stream from another stream, perturbing $\mathcal{D}_{\text{spec}}$ while preserving activation magnitudes. We evaluate on a matched basis: each corrupted evaluation is paired with an uncorrupted baseline using identical samples and resampling indices. Across 4 sub-experiments with different random seeds, we collect N=128 paired samples per task. This paired design maximizes statistical power by controlling sample-level variance, analyzed via paired t-tests with Fisher meta-analysis.
Results. Table 3 provides statistically robust evidence that neural diversity causally affects performance. All three tasks show highly significant accuracy drops ($p<0.001$) when stream-level substitution perturbs diversity ($\Delta\mathcal{D}_{\text{spec}}\approx 0.025$). While effect sizes are modest (0.3% to 0.7% score reduction) — likely because artificial stream substitution creates out-of-distribution corruption patterns — the statistical significance establishes causality beyond correlational association.
Task $\Delta\mathcal{D}_{\text{spec}}$ $\Delta$ Score SE d p-value Sig. N
HaluEval-Summ 0.024 -0.005 0.010 0.007 $1.6\times 10^{-5}$ *** 512
MemoTrap v2 0.031 -0.003 0.010 0.000 $8.2\times 10^{-5}$ *** 512
TruthfulQA-MC2 0.025 -0.007 0.009 0.018 $3.3\times 10^{-7}$ *** 512
Table 3: Artificial corruption of neural diversity establishes statistical causality. Perturbing neural diversity ($\Delta\mathcal{D}_{\text{spec}}>0$) causes accuracy drops across tasks with high statistical significance ($p<0.001$) via paired t-tests with Fisher meta-analysis (N=4 sub-experiments × 128 samples each).
5.2 Ablations
To isolate the contributions of ND-LoRA, we systematically ablate ND-LoRA components at fixed $P=4$ streams. All variants maintain parameter parity through LoRA rank adjustments, enabling fair comparison. We measure run-time spectral diversity ($\mathcal{D}_{\text{spec}}$) at the aggregation layer using evaluation samples, quantifying actual cross-stream correlation during inference.
[…]
5.3 Computational Considerations
[…]
Variant Streams LoRA Regul. Target $\mathcal{D}_{\text{spec}}$ $\overline{\Delta}$% Score $\Delta$ Cost
Standard 1 Single D All 0.0%         1.0x / 1.0x
ParScale $P$ Single D All 0.9991 +0.5% 1.00003x / 1.1x
ParScale-BT $P$ Single D + BT All 0.9988 +1.4% 1.00003x / 1.1x
Stream LoRA $P$ Stream D All 0.2678 +2.9% 1.00003x / 1.1x
Stream LoRA-BT $P$ Stream D + BT All 0.1400 +4.9% 1.00004x / 1.1x
ND-LoRA $P$ Stream D + BT KVQ 0.3755 +12.8% 1.00004x / 1.1x
Table 4: Ablations reveal super-linear combination of impact. Stream LoRA (+2.9%) and Barlow Twins (+1.4%) combine super-linearly (+4.9%), and focusing on KVQ attention amplifies to +12.8%. LoRA: single shared vs. $P$ stream-aware adapters. Regularization: Dropout vs. Barlow Twins. Target: All layers vs. KVQ attention only. $\mathcal{D}_{\text{spec}}$: spectral diversity (lower better). $\overline{\Delta}$% Score: avg. change (hallucination benchmarks).
6 Related Work
Hallucination in Language Models. Hallucinations represent a fundamental challenge in language model deployment (Huang et al., 2024; Tonmoy et al., 2024), proven mathematically inevitable in computable models (Xu et al., 2024; Kalai & Vempala, 2024), with particular severity in smaller models (Lin et al., 2021; Li et al., 2023a). Mechanistic investigations reveal hallucinations arise from internal representation failures (Yu et al., 2024; Ferrando et al., 2025). Current mitigation strategies—retrieval augmentation (Niu et al., 2024), specialized decoding (Li et al., 2023b; Manakul et al., 2023; Wei et al., 2024), and constitutional training (Bai et al., 2022)—address only specific aspects. In other contexts, MoE-LLaVA improves by 1.1% on POPE but requires $8\times$ parameters (Zhang et al., 2024). Our work achieves hallucination reduction through neural diversity within a single architecture at fixed parameter budgets.
Parallel Architectures and Scaling Laws. Parallel scaling offers a third axis beyond parameters and data (Chen et al., 2025; Kaplan et al., 2020). ParScale achieves $O(\log P)$ gains with 22× less memory and 6× less latency (Chen et al., 2025), while inference-optimal scaling laws favor smaller models with more data (Sardana & Frankle, 2024) and MoE achieves 1000× capacity through conditional computation (Shazeer et al., 2017). However, these approaches target accuracy rather than reliability. We show parallel scaling reduces hallucinations through explicit neural diversity.
[…]
Redundancy Reduction and Self-Supervised Learning. Self-supervised methods provide differentiable diversity mechanisms. Barlow Twins prevents collapse through dimension decorrelation (Zbontar et al., 2021), VICReg decomposes into variance, invariance, and covariance terms (Bardes et al., 2022), and understanding dimensional collapse (Jing et al., 2022) reveals why explicit decorrelation is necessary. We adapt these principles for hallucination reduction beyond visual learning.
Parameter-Efficient Fine-Tuning. PEFT methods enable distinct streams under fixed budgets. LoRA reduces parameters by 10,000× with multiple adapters inducing specialization (Hu et al., 2022; Wang et al., 2023), prefix-tuning uses 0.1% of parameters (Li & Liang, 2021), and BitFit shows minimal differences enable specialization when regularized (Ben Zaken et al., 2022). BatchEnsemble and LoRA-Ensemble achieve diversity through clever parameterization (Wen et al., 2020; Mühlematter et al., 2024), providing our technical foundation.
Inference-Time Scaling and Aggregation. Self-consistency improves reasoning by 17.9% through diverse sampling (Wang et al., 2022), though resampling has limitations with imperfect verifiers (Stroebl et al., 2024). Confidence-based weighting reduces required paths by 40% (Taubenfeld et al., 2025), while contrastive decoding and classifier-free guidance achieve improvements equivalent to parameter doubling (Li et al., 2023b; Sanchez et al., 2023). Unlike post-hoc aggregation requiring multiple forward passes, our training-time parallelism learns coordinated streams.
Theoretical Foundations. PAC-Bayesian bounds show minimax-optimal rates connecting diversity to generalization (Steffen et al., 2024; Biggs & Guedj, 2022; Ortega et al., 2022), while concentration inequalities demonstrate that reducing predictor correlation tightens tail bounds on error events (Alquier, 2024). This supports our approach of explicitly regularizing cross-stream correlation for factual reliability.
7 Discussion
Neural diversity establishes a principled framework bridging portfolio theory and deep learning to develop models that are architecturally robust to noise. Our work provides both theoretical grounding — proving hallucination probability is bounded by cross-stream correlation — and empirical validation through ND-LoRA’s 25.6% hallucination reduction at negligible computational cost. The framework reveals why existing scaling approaches target accuracy over reliability and how explicit decorrelation through neural diversity provides a mechanism to reduce hallucinations at fixed parameter budgets, introducing a third scaling axis orthogonal to parameters and data. Unlike traditional ensemble methods (Lakshminarayanan et al., 2017), which require multiple separate models, our approach achieves diversity benefits within a single architecture.
At the same time, our demonstration has limitations. Computational constraints restricted experiments to 0.5B models; larger-scale production-ready implementations are part of planned future work. Benchmark scope focuses on English hallucination and knowledge tasks, with some tasks showing no benefit from increased neural diversity, highlighting task-dependent efficacy that requires further investigation. Finally, while our theoretical framework predicts non-monotonicity, it doesn’t provide a satisfying explanation of why different tasks have different optimality points.
As language models become critical infrastructure, reliability techniques that avoid massive resource investment become essential. While hallucinations may be mathematically inevitable (Kalai & Vempala, 2024), neural diversity demonstrates they can be systematically reduced through architectural design. Our results show reliable AI may emerge from thoughtfully designed architectures rather than brute-force scaling alone, opening new avenues for efficient, trustworthy language models.
Author Contributions
Kushal Chakrabarti led the overall project design, core implementation of ND-LoRA, experimental design, and paper writing. Nirmal Balachundhar contributed theoretical insights and implementation for novel neural diversity techniques, experiment implementation and analysis, and paper writing.
Acknowledgments
We thank the research community at South Park Commons for valuable discussions and feedback throughout this project. We are grateful to Augustine Mavor-Parker, Bhav Ashok, Daniel Morillo, David Sontag, Iman Modarressi, Jaclyn Lunger, John Bohannon, Nelson Ray, Patrick Bozeman and Suhith Rajesh for their insightful comments on earlier drafts. This work was supported by South Park Commons and Modal, who generously provided compute resources for our experiments.
Appendix A Appendix
A.1 Full Proofs
A.1.1 Proof of Lemma 1
Proof.
By bilinearity of covariance,
$\mathrm{Var}(M)$ $=\mathrm{Var}\!\Big(\frac{1}{P}\sum_{i=1}^{P}m_{i}\Big)=\frac{1}{P^{2}}\Bigg(\sum_{i=1}^{P}\mathrm{Var}(m_{i})\;+\;2\sum_{1\leq i<j\leq P}\mathrm{Cov}(m_{i},m_{j})\Bigg)$
$=\frac{1}{P^{2}}\Big(P\,\sigma^{2}\;+\;2\cdot\tbinom{P}{2}\,\rho\,\sigma^{2}\Big)=\frac{1}{P^{2}}\Big(P\,\sigma^{2}+P(P-1)\,\rho\,\sigma^{2}\Big)$
$=\sigma^{2}\!\left(\frac{1}{P}+\frac{P-1}{P}\,\rho\right)=\sigma^{2}\!\left(\frac{1-\rho}{P}+\rho\right).$
A.1.2 Proof of Lemma 2
Proof.
Since $m_{i}=v_{i}^{\top}\tilde{z}_{i}$ and $m_{j}=v_{j}^{\top}\tilde{z}_{j}$,
$\mathrm{Cov}(m_{i},m_{j})=\mathbb{E}\big[(v_{i}^{\top}\tilde{z}_{i})(v_{j}^{\top}\tilde{z}_{j})\big]=v_{i}^{\top}\mathbb{E}[\tilde{z}_{i}\tilde{z}_{j}^{\top}]v_{j}=v_{i}^{\top}C_{ij}v_{j}.$ (14)
By the definition of the spectral norm,
$|\mathrm{Cov}(m_{i},m_{j})|=|v_{i}^{\top}C_{ij}v_{j}|\leq\left\lVert v_{i}\right\rVert\,\left\lVert C_{ij}\right\rVert_{2}\,\left\lVert v_{j}\right\rVert.$ (15)
Divide both sides by $\sigma_{i}\sigma_{j}$ to obtain
$|\rho_{ij}|\leq\frac{\left\lVert v_{i}\right\rVert\,\left\lVert C_{ij}\right\rVert_{2}\,\left\lVert v_{j}\right\rVert}{\sigma_{i}\sigma_{j}}=\kappa_{ij}\,\left\lVert C_{ij}\right\rVert_{2}.$ (16)
Since the right-hand side is nonnegative, this yields the stated upper bound on $\rho_{ij}$. ∎
A.1.3 Proof of Theorem 1
Proof.
Let $\rho_{ij}=\mathrm{Corr}(m_{i},m_{j})$ and $\bar{\rho}\triangleq\frac{2}{P(P-1)}\sum_{i<j}\rho_{ij}$. With common per-stream variance $\sigma^{2}$,
$\mathrm{Var}(M)$ $=\mathrm{Var}\!\Big(\frac{1}{P}\sum_{i=1}^{P}m_{i}\Big)=\frac{1}{P^{2}}\Big(\sum_{i}\mathrm{Var}(m_{i})\;+\;2\sum_{i<j}\mathrm{Cov}(m_{i},m_{j})\Big)$
$=\frac{1}{P^{2}}\Big(P\sigma^{2}+2\textstyle\binom{P}{2}\sigma^{2}\,\bar{\rho}\Big)=\sigma^{2}\!\left(\frac{1}{P}+\frac{P-1}{P}\,\bar{\rho}\right)=\sigma^{2}\!\left(\frac{1-\bar{\rho}}{P}+\bar{\rho}\right).$
Next, by the linear feature model $m_{i}=v_{i}^{\top}\tilde{z}_{i}$ and the definition of $C_{ij}=\mathbb{E}[\tilde{z}_{i}\tilde{z}_{j}^{\top}]$,
$\mathrm{Cov}(m_{i},m_{j})=v_{i}^{\top}C_{ij}v_{j},\qquad|\mathrm{Cov}(m_{i},m_{j})|\leq\|v_{i}\|\,\|C_{ij}\|_{2}\,\|v_{j}\|.$ (17)
Dividing by $\sigma_{i}\sigma_{j}$ gives
$\rho_{ij}\;=\;\frac{\mathrm{Cov}(m_{i},m_{j})}{\sigma_{i}\sigma_{j}}\;\leq\;\frac{\|v_{i}\|\,\|v_{j}\|}{\sigma_{i}\sigma_{j}}\,\|C_{ij}\|_{2}\;=\;\kappa_{ij}\,\|C_{ij}\|_{2}.$ (18)
Averaging over unordered pairs,
$\bar{\rho}=\frac{2}{P(P-1)}\sum_{i<j}\rho_{ij}\;\leq\;\frac{2}{P(P-1)}\sum_{i<j}\kappa_{ij}\,\|C_{ij}\|_{2}=\bar{\kappa}\cdot\frac{2}{P(P-1)}\sum_{i<j}\|C_{ij}\|_{2}=\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}.$
Because $\mathrm{Var}(M)$ is increasing in $\bar{\rho}$ for $P>1$,
$\mathrm{Var}(M)\;\leq\;\sigma^{2}\!\left(\frac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right).$ (19)
Applying Cantelli’s inequality to $M$ with $\mathbb{E}[M]=\mu>0$ and $t=\mu$,
$\mathbb{P}(M\leq 0)=\mathbb{P}(M-\mu\leq-\mu)\leq\frac{\mathrm{Var}(M)}{\mathrm{Var}(M)+\mu^{2}}\leq\frac{\sigma^{2}\left(\tfrac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right)}{\sigma^{2}\left(\tfrac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right)+\mu^{2}}.$ (20)
Finally, if assumptions (ii)–(iii) hold only on a “good” event $G$ with $\mathbb{P}(G)\geq 1-h_{0}$, then
$\mathbb{P}(M\leq 0)=\mathbb{P}(M\leq 0\mid G)\,\mathbb{P}(G)+\mathbb{P}(M\leq 0\mid G^{\mathrm{c}})\,\mathbb{P}(G^{\mathrm{c}})\leq\mathbb{P}(M\leq 0\mid G)+h_{0},$ (21)
and the displayed fraction bounds $\mathbb{P}(M\leq 0\mid G)$, yielding the stated result. ∎
A.1.4 Proof of Theorem 2
Proof.
Let $g(P)\triangleq\tfrac{1-\rho(P)}{P}+\rho(P)$ so that $\mathcal{B}(P)$ is an increasing function of $g(P)$ (monotonicity in the numerator). Compute
$g^{\prime}(P)=\frac{d}{dP}\!\left(\frac{1-\rho(P)}{P}\right)+\rho^{\prime}(P)=\Big(-\frac{\rho^{\prime}(P)}{P}-\frac{1-\rho(P)}{P^{2}}\Big)+\rho^{\prime}(P)=\rho^{\prime}(P)\Big(1-\frac{1}{P}\Big)-\frac{1-\rho(P)}{P^{2}}.$ (22)
As $P\downarrow 1$, we have $1-\tfrac{1}{P}\to 0$ while $1-\rho(P)\to 1-\rho_{0}>0$, hence $g^{\prime}(P)\to-(1-\rho_{0})<0$. For large $P$, the negative term is $O(P^{-2})$, while $\rho^{\prime}(P)=\beta\gamma(P-1)^{\gamma-1}>0$; thus for sufficiently large $P$, $g^{\prime}(P)>0$. By continuity, there exists $P_{\star}>1$ with $g^{\prime}(P_{\star})=0$, implying that $g$ (and therefore $\mathcal{B}$) decreases for $P<P_{\star}$ and increases for $P>P_{\star}$.
If $\gamma\geq 1$, then $\rho^{\prime\prime}(P)=\beta\gamma(\gamma-1)(P-1)^{\gamma-2}\geq 0$ and
$g^{\prime\prime}(P)=\rho^{\prime\prime}(P)\Big(1-\frac{1}{P}\Big)+\frac{2\rho^{\prime}(P)}{P^{2}}+\frac{2(1-\rho(P))}{P^{3}}\;\geq\;0\quad\text{for }P>1.$ (23)
Hence $g$ is convex on $(1,\infty)$ and has a unique minimizer; the same holds for $\mathcal{B}$, which is an increasing transform of $g$. ∎
A.2 Training Cost and Latency Analysis
This appendix provides a complete analysis of computational costs for ND-LoRA and baseline variants. Three key insights enable negligible overhead: (1) fine-tuning on 20M tokens amortizes to less than 0.004% of 1T pretraining, (2) frozen backbone parameters make backward passes nearly free, and (3) inference uses identical FLOPs to ParScale via dynamic LoRA swapping per stream.
A.2.1 Cost Model
Standard Fine-Tuning (P=1) Baseline. Consider a standard LoRA fine-tuning setup with 495M backbone parameters frozen and 1.3M trainable adapter parameters. A typical training step consists of:
  • Forward pass: $1.0\times$ computational cost through 495M parameters
  • Backward pass: $2.0\times$ computational cost through 495M parameters (typical 2:1 backward:forward ratio)
  • Total baseline: 3.0 cost units per training step
ND-LoRA (P=4) Fine-Tuning. With $P=4$ parallel streams, ND-LoRA processes data through multiple independent pathways:
  • Forward pass: $4.0\times$ cost (P parallel forward passes through full 495M model)
  • Backward pass: $2.0\times(1.3\text{M}/495\text{M})\approx 0.005\times$ cost (gradients only propagate through 1.3M trainable parameters after aggregation)
  • Barlow Twins regularization: $1.6\times$ cost (cross-correlation computation across P choose 2 streams and whitening)
  • Prefix/aggregator overhead: $0.05\times$ cost (additional trainable components)
  • Total: 5.655 cost units per training step
Relative Training Cost. The training cost of ND-LoRA relative to standard fine-tuning is:
$\text{Relative cost}=\frac{5.655}{3.0}=1.885\times$ (24)
This is substantially lower than the naive estimate of $4\times$ because backward passes through frozen parameters are essentially free.
A.2.2 Amortization Over Pretraining
To contextualize fine-tuning costs, we amortize over typical pretraining budgets. Given:
  • Pretraining: 1T tokens at $1.0\times$ cost = 1T token-equivalents
  • Fine-tuning: 20M tokens at $1.485\times$ cost = 29.7M token-equivalents
  • Total: $(1\text{T}+37.7\text{M})/1\text{T}=1.0000377\approx 1.00004\times$
The amortized cost is less than 0.004% incremental overhead over the full model lifecycle.
A.2.3 All Variants During Fine-Tuning
Table 5 shows the complete cost breakdown for all ablation variants. The key differences are:
  • Shared vs. Stream-Aware LoRA: Stream-aware adapters add $0.04\times$ prefix overhead
  • Barlow Twins: Adds $1.6\times$ (full BT) or $0.1\times$ (ParScale-BT with simpler correlation)
  • All P=4 variants: Incur $4.0\times$ forward pass cost but only $0.005\times$ backward cost
Variant Forward Backward BT Other Total Relative
Standard 1.0 2.0 0.0 0.0 3.0 1.000×
ParScale 4.0 0.005 0.0 0.01 4.015 1.337×
ParScale-BT 4.0 0.005 0.1 0.01 4.155 1.384×
Indep. LoRA 4.0 0.005 0.0 0.05 4.055 1.352×
ND-LoRA 4.0 0.005 1.6 0.05 5.655 1.885×
Table 5: Fine-tuning cost breakdown (20M tokens). Forward: P parallel passes through 495M backbone. Backward: single pass through 1.3M trainable parameters. BT: Barlow Twins correlation computation. Other: prefix/aggregator overhead. Relative: cost relative to 3.0× standard baseline.
A.2.4 Inference Latency
At inference, all $P>1$ variants exhibit $1.1\times$ latency relative to standard models:
  • Parameter parity: All variants maintain identical total parameter counts by adjusting LoRA rank
  • Parallel processing: P streams process in parallel; latency dominated by slowest stream + aggregation
  • Dynamic loading: Different LoRA adapters are dynamically loaded per stream without duplication
  • Aggregation overhead: Lightweight MLP aggregator adds $\sim$10% latency
The 1.1× factor is consistent across ParScale, ParScale-BT, Indep. LoRA, and ND-LoRA because inference does not involve Barlow Twins regularization and all parameter operations are equivalent.
A.2.5 Summary
  • Training overhead: 1.49× (not 4×) due to free backward passes through frozen parameters
  • Amortized cost: $\leq 0.004\%$ when amortized over 1T-token pretraining
  • Inference latency: 1.1× across all $P\geq 1$ variants with parameter matching
  • Practical impact: Negligible computational overhead for 25.6% hallucination reduction
A.3 LoRA Module Ablations
To understand which components of ND-LoRA contribute most to hallucination reduction, we perform targeted ablations by removing LoRA adapters from specific module types. We compare the full ND-LoRA baseline against two variants:
  • No MLP: removing LoRA from MLP projections (gate_proj, up_proj, down_proj) while keeping attention LoRA
  • No Attention: removing LoRA from attention projections (q_proj, k_proj, v_proj, o_proj) while keeping MLP LoRA
Task $\Delta$% No MLP $\Delta$% No Attention
HaluEval Dialog -1.7% -0.6%
HaluEval QA +16.8% -1.8%
HaluEval Summarization -5.3% -27.0%
MemoTrap v2 +2.5% +0.9%
NQ (8-shot) +11.7% -1.7%
PopQA -0.8% -0.8%
TriviaQA (8-shot) -5.0% -6.9%
TruthfulQA MC1 +3.1% +2.4%
TruthfulQA MC2 +0.2% +1.4%
Table 6: LoRA module ablation results (relative percentage changes from baseline). Evaluations performed on N=1024 samples per task.
A.4 Complete Benchmark Results
Tables 79 provide comprehensive results across $P\in\{1,2,4,8\}$ configurations with parameter-matched $P=1$ baselines. This complete view demonstrates the thoroughness of our evaluation and enables independent verification of claims in the main text.
Evaluation Qwen LoRA ParScale ND-LoRA
HE Dialog 0.458 0.453 0.513
HE QA 0.365 0.337 0.406
HE Summ 0.400 0.439 0.481
MemoTrap 0.634 0.638 0.666
NQ-8 0.065 0.059 0.055
TQA-8 0.188 0.185 0.160
TF-MC1 0.251 0.259 0.269
TF-MC2 0.403 0.412 0.442
NQ-swap 0.550 0.546 0.528
PopQA 0.111 0.109 0.101
Wikitext BPB 0.775 0.797 0.797
Winogrande 0.572 0.564 0.574
Table 7: Benchmark results for R32 (P=2) parameter-matched models.
Evaluation Qwen LoRA ParScale ND-LoRA
HE Dialog 0.464 0.459 0.516
HE QA 0.341 0.322 0.451
HE Summ 0.394 0.409 0.502
MemoTrap 0.629 0.634 0.635
NQ-8 0.065 0.061 0.059
TQA-8 0.191 0.185 0.172
TF-MC1 0.245 0.253 0.262
TF-MC2 0.399 0.413 0.416
NQ-swap 0.554 0.542 0.535
PopQA 0.110 0.110 0.106
Wikitext BPB 0.778 0.793 0.795
Winogrande 0.564 0.573 0.577
Table 8: Benchmark results for R64 (P=4) parameter-matched models.
Evaluation Qwen LoRA ParScale ND-LoRA
HE Dialog 0.460 0.465 0.475
HE QA 0.344 0.335 0.370
HE Summ 0.379 0.416 0.450
MemoTrap 0.630 0.639 0.689
NQ-8 0.066 0.063 0.059
TQA-8 0.192 0.182 0.171
TF-MC1 0.251 0.256 0.259
TF-MC2 0.407 0.414 0.424
NQ-swap 0.551 0.540 0.554
PopQA 0.110 0.109 0.103
Wikitext BPB 0.778 0.779 0.784
Winogrande 0.569 0.577 0.568
Table 9: Benchmark results for R128 (P=8) parameter-matched models.
A.5 Use of Large Language Models
Large language models were used as a compilation tool to assist with writing and organizing sections of this paper, including literature review synthesis, section structuring, LaTeX formatting, and co-generation of experimental code. All technical content, experimental design, theoretical contributions, and scientific claims are the authors’ original work. The models served primarily to improve clarity, organization, and implementation of our ideas rather than generate novel scientific insights.